# Triad: A Framework Leveraging a Multi-Role LLM-based Agent to Solve Knowledge Base Question Answering

### v3

@@ Abstract
Chang Zong Zhejiang University zongchang@zju.edu.cn                        Yuchen Yan Zhejiang University yanyuchen@zju.edu.cn                        Weiming Lu Zhejiang University luwm@zju.edu.cn
Eliot Huang Leibowitz AI EliotHuang@leibowitz.org.cn                        Jian Shao Zhejiang University jshao@zju.edu.cn                        Yueting Zhuang Zhejiang University yzhuang@zju.edu.cn
@@ 1 Introduction
Studies on KBQA with LLMs has attracted considerable attention. Some works focus primarily on highlighting the inability of LLMs to generate complete factoid results Hu et al. (2023b); Tan et al. (2023c) or demonstrating their potential efficacy in future research Omar et al. (2023b); Tan et al. (2023b). Other works concentrates on generating answers by prompt learning and incorporating external knowledge bases Baek et al. (2023); Tan et al. (2023a). Concurrently, LLMs can be deployed to address each phase within Text2SQL challengesLi et al. (2023, 2024) or theorem proof tasksDong et al. (2023). However, each phase of KBQA can be further decomposed into subtasks and completed through an agentic approach that provides feedback and cooperation. Additionally, decomposing the task reduces the complexity of cooperative working by allowing each role to concentrate on smaller sub-problemsWang et al. (2020). As illustrated in Figure 1, three roles in an organization work together to provide the final answer for the overall task. The above observations spur our exploration into the following question: Whether an LLM-based agent can solve KBQA tasks by serving as multiple roles.
@@ 5.2 LLM-based Agents for Complex Tasks
LLMs have recently gained significant attention due to their ability to approximate human-level intelligence through the acquisition of vast amounts of web knowledge. This has led to numerous studies focusing on LLM-based agents. A recent surveyWang et al. (2023) proposes a unified architectural design for LLM-based agents, which consists of four modules that include profile, memory, plan, and action. CHATDBHu et al. (2023a) employs an LLM controller to generate SQL instructions, enabling it to manipulate databases, which allows for symbolic memory and complex multi-hop reasoning. ARTParanjape et al. (2023) uses a frozen LLM to generate intermediate reasoning steps and further integrates tools for new tasks with minimal human intervention. ToolformerSchick et al. (2024) takes a different approach by training an LLM to plan and execute tools for the next token prediction by learning API calls generation. ReActYao et al. (2023) focuses on overcoming LLM hallucination by interacting with external knowledge bases, thus generating interpretable task-solving strategies. Divergent from the aforementioned studies, our framework concentrates on the solving KBQA tasks by introducing a multi-role LLM-based agent that specializes in various subtasks distributed across different phases.
@@ 6 Conclusion
Appendix A Prompts Provided to LLMs of G-Agent for Solving Various Subtasks in KBQA
@@
Appendix B Prompts Provided to LLMs of D-Agent for Solving Selection Subtasks in KBQA
@@
Appendix C Prompt Provided to LLMs of A-Agent for Solving Answering Subtask in KBQA

### v4

@@ Abstract
Chang Zong Zhejiang University zongchang@zju.edu.cn                        Yuchen Yan Zhejiang University yanyuchen@zju.edu.cn                        Weiming Lu Zhejiang University luwm@zju.edu.cn                        Jian Shao Zhejiang University jshao@zju.edu.cn
Eliot Huang Leibowitz AI EliotHuang@leibowitz.org.cn                        Heng Chang Tsinghua University changh17@tsinghua.org.cn                        Yueting Zhuang Zhejiang University yzhuang@zju.edu.cn
@@ 1 Introduction
Studies on KBQA with LLMs has attracted considerable attention. Some works focus primarily on highlighting the inability of LLMs to generate complete factoid results Hu et al. (2023b); Tan et al. (2023c) or demonstrating their potential efficacy in future research Omar et al. (2023b); Tan et al. (2023b). Other works concentrates on generating answers by prompt learning and incorporating external knowledge bases Baek et al. (2023); Tan et al. (2023a). Concurrently, LLMs can be deployed to address each phase within Text2SQL challengesLi et al. (2023, 2024) or theorem proof tasksDong et al. (2023). However, each phase of KBQA can be further decomposed into subtasks and completed through an agentic approach that provides feedback and cooperation. Additionally, decomposing the task reduces the complexity of cooperative working by allowing each role to concentrate on smaller sub-problemsWang et al. (2020). As illustrated in Figure 1, three roles in an organization work together to provide the final answer for the overall task. The above observations spur our exploration into the following question: How does an LLM-based agent solve KBQA tasks by serving as multiple roles, and its performance is comparable to systems trained specifically?
@@ 5.2 LLM-based Agents for Complex Tasks
LLMs have recently gained significant attention due to their ability to approximate human-level intelligence. This has led to numerous studies focusing on LLM-based agents. A recent surveyWang et al. (2023) proposes a unified architecture for LLM-based agents, which consists of four modules that include profile, memory, plan, and action. CHATDBHu et al. (2023a) employs an LLM controller to generate SQL instructions, which allows for symbolic memory and complex multi-hop reasoning. ARTParanjape et al. (2023) uses a frozen LLM to generate reasoning steps and further integrates tools for new tasks with minimal human intervention. ToolformerSchick et al. (2024) takes a different approach by training an LLM to plan and execute tools for the next token prediction by learning API calls generation. ReActYao et al. (2023) focuses on overcoming LLM hallucination by interacting with external knowledge bases, thus generating interpretable task-solving strategies. CodeAgentTang et al. (2024) designs a multi-agent collaboration system across four phases in a code review process. Divergent from the aforementioned studies, our framework concentrates on the solving KBQA tasks by introducing a multi-role LLM-based agent that specializes in various subtasks distributed across different phases.
@@ 6 Conclusion
Limitations
The limitation of our research lies in a further research and discussion on a more diverse range of multi-agent collaboration manners for KBQA tasks. Moreover, this framework worth evaluating on domain specific QA datasets.
@@
Appendix A Response Time Analysis
We analyze various QA frameworks in response time to a question. The average latency of each phase including question parsing (QP), URI linking (UL), and answer generation (AG) for each knowledge base is reported. We randomly select 10 samples from each dataset for evaluation to obtain the average response times for Triad-1 and Triad-3, which represent retrying three times and generating an answer in one go, respectively, during the answer generation phase. The comparison between traditional QA systems and Triad is shown in Figure 3. Triad generally shows a competitive time-consuming performance to latest traditional QA systems. Specifically, compared to other phases, URL linking consumes more time due to the need to invoke LLM multiple times. Moreover, according to Section 4.4, with smaller retry times of A-Agent, Triad can significantly reduce time cost while only causing slight performance degradation, revealing the advantages of our framework in balancing performance and efficiency.
Figure 3: Response time of traditional KBQA systems and Triad on three datasets. Each bar shows average response time of a particular phase of KBQA.
Appendix B Prompts Provided to LLMs of G-Agent for Solving Various Subtasks in KBQA
@@
Appendix C Prompts Provided to LLMs of D-Agent for Solving Selection Subtasks in KBQA
@@
Appendix D Prompts Provided to LLMs of A-Agent for Solving Answering Subtask in KBQA